The sum of the ages of Mark and Linda is 114 years. 7 years ago, Mark's age was 4 times Linda's age. How old is Mark now?

Answer by JulietG(1812)

M + L = 114
M-7 = 4(L-7)
Distribute
M-7 = 4L-28
Add 7 to each side
M = 4L-21
Take that value and substitute it into the first equation
M + L = 114
(4L-21)+L = 114
5L-21 = 114
Add 21 to each side
5L = 135
Divide each side by 5
L = 27
If Linda is 27, then Mark is 114-27, or 87.

Prove it:
7 years ago, Mark's age was 4 times Linda's age
7 years ago, Mark would have been 87-7, or 80 years old
7 years ago, Linda would have been 27-7, or 20 years old
80 years old is 4 times as old as 20.